# Generating Jackiw-Teitelboim Euclidean gravity from static three-dimensional Maxwell-Chern-Simons electromagnetism

### v1

@@ Abstract
We consider pure three-dimensional Maxwell-Chern-Simons electrodynamics in the static limit. We show that this theory can be mapped onto a two-dimensional gravitational model in the first-order formalism of Riemannian manifolds with Euclidean signature, coupled to a real scalar field naturally interpreted as a dilaton. In this framework, the Newtonian and cosmological constants in two dimensions are fully determined by the electric charge. The solution to this gravitational model is found to be trivial: a constant dilaton field on a flat manifold. However, we introduce two distinct shifts of the spin-connection that transform the model into Jackiw-Teitelboim gravity. Specifically, we identify two additional solutions: a hyperbolic manifold with also a constant dilaton configuration; and a spherical manifold where, again, the dilaton assumes a constant, nonzero field configuration. In both non-flat cases, by employing the Gauss-Bonnet theorem in the specific cases of compact manifolds, we establish that the manifold’s radius is fixed by the cosmological constant (and, therefore, by the electric charge).
@@ 1 Introduction
The primary objective of this work is to establish a connection between three-dimensional electrodynamics and Jackiw-Teitelboim (JT) gravity models [1, 2, 3, 4, 5]. Accordingly, the main system of interest in this paper is two-dimensional gravity [1, 2, 6, 7, 8, 9, 10, 11, 3, 5, 12]. In particular, two-dimensional gravity has served as a robust framework for probing the quantization of gravity in higher dimensions. Its topological nature further motivates the investigation of topological field theories as candidates for quantum gravity in four dimensions (See, for instance, [8, 13, 14, 15] and references therein). Another compelling reason for studying gravity theories across different dimensions is that numerous physical systems can be effectively described by geometrodynamical models [16, 17, 18, 19, 20, 21, 22]. In such models, gravity (or geometrodynamics) provides an alternative framework for describing certain gauge theories.
@@
Specifically, in the case of JT gravity, several potential applications and relevant results have been explored [23, 24, 25, 26, 27, 28, 4, 29, 5, 30, 31, 32]. It is particularly significant in black hole physics. While the Einstein-Hilbert model with cosmological constant in two dimensions simply defines a Plateau problem [33, 34], the JT model has more dynamical features due to the presence of the dilaton, being thus a good description for two-dimensional black holes and for a large group of near-extremal black holes in higher dimensions [5, 24, 35, 26]. Another area of great contribution from JT gravity is holography (see [36] and references therein). Typically, JT models have constant and negative curvature as solutions, stating an $AdS_{2}$ spacetime (two-dimensional Anti-de Sitter spacetime). In this manner, $AdS/CFT$ correspondence [37] can play an important role in this scenario in order to study quantum gravity in a string theory fashion. By analyzing such a lower-dimensional system, the structure of quantum gravity could emerge more evidently and provide some insights for higher dimensions. In addition, JT gravity is relevant in matrix models [38, 39], which can be related to Sachdev–Ye–Kitaev (SYK) systems [40, 41] to describe chaotic quantum systems in condensed matter physics. It turns out that the SYK model is deeply connected to black hole physics due to a duality to JT gravity via $AdS/CFT$ correspondence (see for example [42] and references therein).
@@
The second system under consideration in this work is three-dimensional electrodynamics, which allows for the inclusion of the Chern-Simons (CS) topological term. Electromagnetic models of this type have applications in various branches of physics, extending beyond purely field theoretical investigations. Notably, they have been employed to describe aspects of topological insulators [43, 44], Weyl semi-metals and axion effects [45, 46], graphene layers [47, 48, 49, 50], and even three-dimensional gravity [51, 16, 52, 21, 22]. Furthermore, the consistency of CS theories at the quantum level is a highly valued feature. In fact, it has long been established that quantum CS theories are perturbatively finite [53, 54].
@@
The approach we adopt to connect electrodynamics to gravity theories was previously investigated in two recent studies by the authors [55, 56]. In those works, we demonstrated that two-dimensional gravity can emerge from two-dimensional electrodynamics (including fermions) by defining a suitable mapping between the fields of both theories. This mapping originates from the isomorphism between the groups $U(1)$ and $SO(2)$. It is important to highlight that the mapping is rigorously defined in Euclidean spaces — i.e., electrodynamics formulated in $\mathbb{R}^{2}$ spacetime is mapped onto gravity in a two-dimensional manifold equipped with a Euclidean signature metric. As will become clear throughout the paper, this feature is highly beneficial in the present study.
@@
At this point, it is also worth mentioning that the gravity models examined here are formulated within the first-order formalism (FOF) [57, 58, 59, 33, 34]. The FOF is a theoretical framework in which gauge symmetry is identified with local spacetime isometries, bringing gravity more closely related to standard gauge theories. Moreover, in this formalism, the fundamental fields are the vielbein (zweibein, in the specific case of two dimensions) and the spin-connection, rather than the metric tensor and affine connection used in the Palatini formalism [60], although the correspondence between both formalisms is well defined [61].
@@
Furthermore, we identify a Euclidean two-dimensional dilatonic gravity model whose solution is trivial: the dilaton field remains constant throughout the manifold, while the manifold itself is flat with minimal area. Nonetheless, we demonstrate that the model can be reformulated to describe JT gravity with two distinct nontrivial solutions. The first corresponds to a manifold with a negative constant curvature and a constant dilaton, i.e., the model admits hyperbolic spaces as solutions [63, 64]. The second non-flat solution describes a regular spherical manifold with, again, a constant dilaton configuration.
In the three cases (flat, hyperbolic, and spherical), the Gauss-Bonnet (GB) theorem can be employed to determine some extra properties of these manifolds. Particularly, if we consider compact surfaces, we can study: the flat torus; the $n$-holed torus (negative curvature); and the sphere (positive curvature). Except for the flat case, we are able to determine the length typical scale of such surfaces, finding it to be of the order of $\mu m$ (micrometers). This is possible since the GB theorem relates the area of surfaces to their Euler characteristics [65, 63], and because the area depends on the cosmological constant (explicitly calculated from the electric charge). In addition, it is crucial to emphasize that the CS term plays a fundamental role in achieving the JT model. In fact, within our framework, the JT term is directly proportional to the CS level of the theory.
This work is organized as follows. In Section 2, we define three-dimensional MCS electrodynamics, discuss its main features, and take its static limit. In Section 3 we map the static electrodynamics into a two-dimensional gravity theory in the FOF and explore its flat solution. Then, in Section 4 we connect this gravity theory with JT gravities and explore the main properties of these solutions. In Section 5, we compute, by topological arguments, the radius of some manifolds, specifically for the $n$-holed torii and the sphere. Finally, our conclusions and perspectives are displayed in Section 6.
@@ 2.1 The Maxwell-Chern-Simons theory
The starting point is the most general action in three dimensions of pure electrodynamics which is polynomial in the fields and their derivatives and power counting renormalizable. The result is just the pure MCS action,
$S_{3qed}=\int\left(-\frac{1}{e^{2}}f\underline{\ast}f+\kappa af\right)\;,$ (2.1)
with $a=a_{\mu}dx^{\mu}$ being the gauge field, $f=\underline{d}a$ being the field strength while $\underline{\ast}$ characterizes the three-dimensional Hodge dual. The notation $\underline{d}=\partial_{\mu}dx^{\mu}$ stands for the three-dimensional exterior derivative. Greek indices run through $\{0,1,2\}$ with the 00th coordinate characterizing time. Moreover, we consider Minkowskian three-dimensional spacetime with negative signature as the stage of the MCS theory (2.1). We also remark that we are using natural units.
The first term in (2.1) is clearly the Maxwell action. Therefore, $e$ is the electric charge. The second term in (2.1) is the CS topological term and $\kappa$ is the CS level of the theory. The canonical dimensions of fields and parameters are given by $\left[a\right]=1$, $\left[e\right]=1/2$, and $\left[\kappa\right]=0$. The field equations can easily be derived from the action (2.1) by applying Hamilton’s principle to it. It provides
$\underline{d}\underline{\ast}f+\kappa e^{2}f=0\;,$ (2.2)
Moreover, Bianchi identities are also valid, namely,
@@
By performing a suitable and straightforward spacetime decomposition, one can easily infer that the field strength $f$ is related to the electromagnetic fields through
@@
Such field strength describes an electrodynamics governed by a two-dimensional electric vector field and a pseudo-scalar magnetic field, as expected. The equations of motion can be decoupled in the usual way to generate massive wave equations,
@@
with $\Box=\partial_{t}^{2}-\nabla^{2}$ being the d’Alambertian operator, ${\bf E}=(E^{1},E^{2})$, and $m=\kappa e^{2}$ being the CS topological mass. At the classical level, these waves decay exponentially on a characteristic length scale of the order of $m^{-1}$. This is the usual result of MCS electromagnetism.
Finally, it is convenient to recall that electrodynamics is known to be a gauge theory invariant under $U(1)$ gauge transformations,
@@
with $\alpha$ being a local parameter. In fact, electrodynamics can be fundamentally defined as a $U(1)$ gauge theory [66]. It turns out that the $U(1)$ symmetry is the key for the mapping connecting this theory to gravity.
@@ 2.2 Dimensional reduction - The static limit
Once the MCS is defined, dimensional reduction of the model can be performed. We choose to reduce the time coordinate by considering only static systems. This is the simplest way to reduce the time dimension. Therefore, decomposing space and time of the exterior derivative, the gauge field, and field strength, we find
@@
¯$\underline{\ast}f$ $=$ $-\ast\left(d\Phi-\partial_{t}A\right)+\ast Fdt\;,$ (2.7)
with $A=A_{i}dx^{i}$ being the two-dimensional gauge field, $\Phi$ being the electric potential, and $F=dA$ being the two-dimensional electromagnetic field strength. Moreover, $\ast$ is the Hodge dual in the two-dimensional Euclidean space while $d=\partial_{i}dx^{i}$ is the purely spatial exterior derivative. The Latin indices $i,j,k\ldots$ run through $\{1,2\}$, and describe the Euclidean spatial sector. Thus, the Lagrangian terms decompose accordingly,
@@
[…]
$S_{3stqed}=\int\left[-\frac{1}{e^{2}}\left(d\Phi\ast d\Phi+F\ast F\right)+\kappa\left(\Phi F+Ad\Phi\right)\right]\;,$ (2.10)
where integration is taken over the two-dimensional Euclidean space. Moreover, since $Ad\Phi=\Phi F-d(\Phi A)$, we are allowed to rewrite action (2.10) as
$S_{3stqed}=\int\left\{-\frac{1}{e^{2}}\left(d\Phi\ast d\Phi+F\ast F\right)+\kappa\left[2\Phi F-d\left(\Phi A\right)\right]\right\}\;,$ (2.11)
simplifying thus the term $Ad\Phi$ to a surface term. Of course, the surface term can be relevant in systems with non-trivial boundaries, like topological insulators or similar systems.
The action (2.11) is just a sort of three-dimensional static electromagnetism that can be viewed in different ways. For instance, it can be interpreted as electrodynamics in two Euclidean dimensions coupled to a scalar field; equivalently, still under the two-dimensional point of view, one can interpret the static model as a scalar field $\Phi$ interacting with a pseudo-scalar magnetic field through the reduced field strength
@@
It is worth mentioning that the present situation is essentially different from the previous works [55, 56] where the two-dimensional electrodynamics considered was purely electrical. It is possible to achieve a purely electrical system by reducing the third dimension ($\mu=2$) instead of the time coordinate. However, the resulting electrical model would still be living in Minkowskian spacetime. Such a model could not be mapped in a gravity model in an easy way as in the Euclidean case. This is the main reason why we dimensionally reduce time instead of a spatial dimension. See [55, 56] for further details.
The static field equations can be derived directly from the field equations (2.2) or from the minimization of the action (2.11). The field equations of the scalar and vector fields read
@@
with $G$ and $C=C_{i}dx^{i}$ being integration constants that can be chosen to vanish, with no loss of generality. These are massive Laplace equations whose solutions, in polar coordinates, are given in terms of the modified Bessel functions of the first and second kinds depending on $mr$ (with $r$ being the radial coordinate), and periodic functions in the angular coordinate.
It is also useful, for future purposes, to infer that, after the dimensional reduction, the gauge transformations (2.6) decouple to
@@
i.e., the scalar field becomes invariant under gauge transformations while the magnetic vector potential remains as a gauge field.
As an extra and final step, we rewrite the action (2.11) with the help of an auxiliary field (See [17, 55, 56]),
$S_{3qed}=\int\left\{-\frac{1}{e^{2}}\left(d\Phi\ast d\Phi+\Theta\ast F-\frac{1}{2}\Theta\ast\Theta\right)+\kappa\left[2\Phi F-d\left(\Phi A\right)\right]\right\}\;.$ (2.16)
The field equation of the auxiliary field $\Theta$ trivially reads $\Theta=F$, which implies in the action (2.11) again. Of course, the introduction of auxiliary fields in this way is also valid at the quantum level [67].
To end up this Section, for completeness, and usefulness, the canonical dimensions, and form ranks of the fields and parameters appearing in the action (2.16) are displayed in Table 1.
Fields $A$ $\Theta$ $\Phi$ $e$ $\kappa$
Dimension $1$ $2$ $1$ $1/2$ 00
Rank $1$ $2$ 00 00 00

### v2

@@ Abstract
We consider pure three-dimensional Maxwell-Chern-Simons electrodynamics in the static limit. We show that this theory can be mapped onto a two-dimensional gravitational model in the first-order formalism of Riemannian manifolds with Euclidean signature, coupled to a real scalar field naturally interpreted as a dilaton. In this framework, the Newtonian and cosmological constants in two dimensions are fully determined by the electric charge. The solution to this gravitational model is found to be trivial: a constant dilaton field on a flat manifold. However, we introduce two distinct shifts of the spin connection that transform the model into Jackiw-Teitelboim gravity. Specifically, we identify two additional solutions: a hyperbolic manifold with also a constant dilaton configuration; and a spherical manifold where, again, the dilaton assumes a constant, nonzero field configuration. In both non-flat cases, by employing the Gauss-Bonnet theorem in the specific cases of compact manifolds, we establish that the manifold’s radius is fixed by the cosmological constant (and, therefore, by the electric charge).
@@ 1 Introduction
The primary objective of this work is to establish a connection between three-dimensional electrodynamics and Jackiw-Teitelboim (JT) gravity models [1, 2, 3, 4, 5, 6], thereby contributing to the broader program of gauge/gravity correspondences. The well-known $AdS/CFT$ duality from string theory [7] asserts that gravitational theories can often be described equivalently by gauge theories. The theories considered in this work are both simple and explicitly solvable, providing a concrete and accessible model through which this duality can be tested and understood in detail, without relying on the full machinery of string theory. Accordingly, the main system of interest in this paper is two-dimensional gravity [1, 2, 8, 9, 10, 11, 12, 13, 3, 6, 14]. In particular, two-dimensional gravity has served as a robust framework for probing the quantization of gravity in higher dimensions. Its topological character further motivates the investigation of topological field theories as candidates for quantum gravity in four dimensions (see, for instance, [10, 15, 16, 17] and references therein). Another compelling reason for studying gravity theories across different dimensions is that numerous physical systems can be effectively described by geometrodynamical models [18, 19, 20, 21, 22, 23, 24]. In such models, gravity (or geometrodynamics) provides an alternative framework for describing certain gauge theories.
@@
Specifically, in the case of JT gravity, numerous potential applications and significant results have been investigated [25, 26, 27, 28, 29, 30, 4, 31, 5, 6, 32, 33, 34]. It plays a particularly important role in black hole physics. While the Einstein-Hilbert model with cosmological constant in two dimensions simply defines a Plateau problem [35, 36], the JT model exhibits richer dynamical features due to the presence of the dilaton, making it a good description for two-dimensional black holes and for a large group of near-extremal black holes close to their horizons in higher dimensions [6, 26, 37, 28, 5]. Thus, establishing a correspondence with Maxwell–Chern–Simons electrodynamics suggests that thermodynamic and quantum features of black holes — such as entropy and near-extremal dynamics — might have analogues in simple electromagnetic gauge systems. Another major area where JT gravity has had a strong impact is holography (see [38] and references therein). Typically, JT models have constant and negative curvature as solutions, stating an $AdS_{2}$ spacetime (two-dimensional Anti-de Sitter spacetime). In this manner, $AdS/CFT$ correspondence [7] plays a central role in studying quantum gravity from a string-theoretic perspective. By analyzing such a lower-dimensional system, the underlying structure of quantum gravity may emerge more transparently, providing valuable insights for higher dimensions. In addition, JT gravity is also relevant in matrix models [39, 40], which are closely related to Sachdev–Ye–Kitaev (SYK) systems [41, 42] to describe chaotic quantum systems in condensed matter physics. It turns out that the SYK model is deeply connected to black hole physics due to a duality to JT gravity via $AdS/CFT$ correspondence (see, for example, [43, 5] and references therein).
@@
The second system considered in this work is three-dimensional electrodynamics, which allows for the inclusion of the Chern-Simons (CS) topological term. Electromagnetic models of this kind have applications in various branches of physics, extending beyond purely field-theoretical investigations. Notably, they have been employed to describe aspects of topological insulators [44, 45], Weyl semimetals and axion effects [46, 47], graphene layers [48, 49, 50, 51], and even three-dimensional gravity [52, 18, 53, 23, 24]. Moreover, the consistency of CS theories at the quantum level is a highly valued feature. In fact, quantum CS theories have long been known to be perturbatively finite [54, 55]. Relating this theory to JT gravity thus enables a reinterpretation of certain topological or confinement features of the gauge system in purely geometric terms, and, conversely, allows field-theoretical tools to be applied to gravitational problems. Additionally, the existence of a mapping between these theories provides an exactly solvable framework to test ideas about gauge/gravity duality, offers a fresh perspective on the quantization of Chern–Simons theories, and opens the possibility of employing condensed matter systems as real laboratory simulators of gravitational dynamics.
@@
The approach we adopt to relate electrodynamics to gravity theories was previously explored in two recent studies by the authors [56, 57]. In those works, we showed that two-dimensional gravity can emerge from two-dimensional electrodynamics (including fermions) by defining a suitable mapping between the fields of both theories. This mapping arises from the isomorphism between the groups $U(1)$ and $SO(2)$. It is important to emphasize that the mapping is rigorously defined in Euclidean spaces — i.e., electrodynamics formulated on $\mathbb{R}^{2}$ spacetime is mapped onto gravity on a two-dimensional manifold equipped with a Euclidean signature metric. As will become evident throughout the paper, this feature is particularly advantageous in the present study.
@@
At this point, it is also worth mentioning that the gravity models considered here are formulated within the first-order formalism (FOF) [58, 59, 60, 35, 36]. The FOF is a theoretical framework in which gauge symmetry is identified with local spacetime isometries, thereby connecting gravity more closely to standard gauge theories. Moreover, in this formalism, the fundamental fields are the vielbein (or zweibein in two dimensions) and the spin connection, rather than the metric tensor and affine connection used in the Palatini formalism [61], although the correspondence between the two formalisms is well defined [62]. Additionally, we assume metric compatibility and vanishing torsion. Note that in the FOF of gravity the gauge curvature is also the spacetime curvature. On the other hand, electrodynamics with the CS term is a gauge theory directly tied to topological invariants, but with a gauge curvature that has no relation to spacetime curvature since its gauge symmetry is an internal one. Establishing a correspondence between both theories would enable one to rephrase topological data in a gauge theory of internal symmetry straightforwardly as geometry, thus bridging topological and geometric approaches.
@@
Furthermore, we identify a two-dimensional Euclidean dilatonic gravity model whose solution is trivial: the dilaton field remains constant throughout the manifold, while the manifold itself is flat with minimal area. Nevertheless, we demonstrate that the model can be reformulated to describe JT gravity with two distinct nontrivial solutions. The first corresponds to a manifold with a negative constant curvature and a constant dilaton, i.e., the model admits hyperbolic spaces as solutions [64, 65]. The second non-flat solution describes a regular spherical manifold again with a constant dilaton configuration.
In the three cases (flat, hyperbolic, and spherical), the Gauss-Bonnet (GB) theorem can be employed to determine additional properties of these manifolds. Particularly, for compact surfaces, we can study: the flat torus; the $n$-holed torus (negative curvature); and the sphere (positive curvature). Except for the flat case, we are able to determine the characteristic length scale of such surfaces, which is found to be on the order of $\mu m$ (micrometers). This is possible because the GB theorem relates the area of surfaces to their Euler characteristics [66, 64], and because the area depends on the cosmological constant (explicitly calculated from the electric charge). In addition, it is crucial to emphasize that the CS term plays a fundamental role in achieving the JT model. In fact, within our framework, the JT term is directly proportional to the CS level of the theory.
This work is organized as follows. In Section 2, we introduce three-dimensional MCS electrodynamics, discuss its main features, and consider its static limit. In Section 3 we map the static electrodynamics onto a two-dimensional gravity theory in the FOF and analyze its flat solution. Then, in Section 4 we relate this gravity theory to JT gravities and examine the main properties of these solutions. In Section 5, we compute, using topological arguments, the radius of some manifolds, specifically for the $n$-holed tori and the sphere. Finally, our conclusions and perspectives are presented in Section 6.
@@ 2.1 The Maxwell-Chern-Simons theory
The starting point is the most general three-dimensional action for pure electrodynamics that is polynomial in the fields and their derivatives and is power-counting renormalizable. The resulting action is simply the pure MCS action,
$S_{3qed}=\int\left(-\frac{1}{\mathrm{e}^{2}}f\underline{\ast}f+\kappa af\right)\;,$ (2.1)
where $a=a_{\mu}dx^{\mu}$ denotes the gauge field, $f=\underline{d}a$ denotes the field strength, and $\underline{\ast}$ represents the three-dimensional Hodge dual. The notation $\underline{d}=\partial_{\mu}dx^{\mu}$ stands for the three-dimensional exterior derivative. Greek indices run over $\{0,1,2\}$ with the 0th coordinate characterizing time. Moreover, we consider a three-dimensional Minkowski spacetime with a negative signature as the background for MCS theory (2.1). We also note that natural units are used throughout.
The first term in (2.1) is clearly the Maxwell action; therefore, $\mathrm{e}$ is the electric charge. The second term in (2.1) is the CS topological term and $\kappa$ denotes the CS level of the theory. The canonical dimensions of fields and parameters are given by $\left[a\right]=1$, $\left[\mathrm{e}\right]=1/2$, and $\left[\kappa\right]=0$. The field equations can be derived straightforwardly from the action (2.1) by applying Hamilton’s principle, yielding
$\underline{d}\underline{\ast}f+\kappa\mathrm{e}^{2}f=0\;,$ (2.2)
Moreover, the Bianchi identities also hold, namely,
@@
By performing a suitable and straightforward spacetime decomposition, one can easily infer that the field strength $f$ is related to the electromagnetic fields as
@@
This field strength describes an electrodynamics governed by a two-dimensional electric vector field and a pseudoscalar magnetic field, as expected. The equations of motion can be decoupled in the standard way to obtain massive wave equations,
@@
with $\Box=\partial_{t}^{2}-\nabla^{2}$ denoting the d’Alembertian operator, ${\bf E}=(E^{1},E^{2})$, and $m=\kappa\mathrm{e}^{2}$ the CS topological mass. At the classical level, these waves decay exponentially over a characteristic length scale of order $m^{-1}$. This is the standard result of MCS electromagnetism.
Finally, it is worth recalling that electrodynamics is known to be a gauge theory invariant under $U(1)$ gauge transformations,
@@
with $\alpha$ being a local parameter. In fact, electrodynamics can be fundamentally formulated as a $U(1)$ gauge theory [67]. It turns out that the $U(1)$ symmetry is the key for the mapping connecting this theory to gravity.
@@ 2.2 Dimensional reduction - The static limit
Once the MCS theory is defined, dimensional reduction of the model can be performed. We choose to reduce the time coordinate by considering only static systems, which provides the simplest way to eliminate the time dimension. Therefore, decomposing space and time of the exterior derivative, the gauge field, and field strength, we obtain
@@
$\underline{\ast}f$ $=$ $-\ast\left(d\Phi-\partial_{t}A\right)+\ast Fdt\;,$ (2.7)
with $A=A_{i}dx^{i}$ denotes the two-dimensional gauge field, $\Phi$ denotes the electric potential, and $F=dA$ is the two-dimensional electromagnetic field strength. Here, $\ast$ is the Hodge dual in the two-dimensional Euclidean space while $d=\partial_{i}dx^{i}$ is the purely spatial exterior derivative. The Latin indices $i,j,k\ldots$ run over $\{1,2\}$, and describe the Euclidean spatial sector. Accordingly, the Lagrangian terms decompose as
@@
$S_{3qed}=\int\left\{-\frac{1}{\mathrm{e}^{2}}\left[\left(d\Phi-\partial_{t}A\right)\ast\left(d\Phi-\partial^{t}A\right)+F\ast F\right]+\kappa\left[\Phi F+A\left(d\Phi-\partial_{t}A\right)\right]\right\}dt\;.$ (2.9)
Dimensional reduction can be carried out by restricting attention to static systems. In this case, all time derivatives vanish, while the factor $\int dt$ appears only as a normalization constant to be removed. The action (2.9) then reduces to the static electrodynamical action
$S_{3stqed}=\int\left[-\frac{1}{\mathrm{e}^{2}}\left(d\Phi\ast d\Phi+F\ast F\right)+\kappa\left(\Phi F+Ad\Phi\right)\right]\;,$ (2.10)
where integration is taken over two-dimensional Euclidean space. Moreover, since $Ad\Phi=\Phi F-d(\Phi A)$, we may rewrite action (2.10) as
$S_{3stqed}=\int\left\{-\frac{1}{\mathrm{e}^{2}}\left(d\Phi\ast d\Phi+F\ast F\right)+\kappa\left[2\Phi F-d\left(\Phi A\right)\right]\right\}\;,$ (2.11)
so that the term $Ad\Phi$ reduces to a surface contribution. This boundary term can, of course, play a role in systems with non-trivial boundaries, such as topological insulators or related setups.
The action (2.11) can be regarded as a form of three-dimensional static electromagnetism, which admits different interpretations. For example, it can be viewed as electrodynamics in two Euclidean dimensions coupled to a scalar field. Equivalently, from the two-dimensional perspective, the static model may be seen as a scalar field $\Phi$ interacting with a pseudo-scalar magnetic field through the reduced field strength
@@
It is worth noting that the present situation differs essentially from the previous works [56, 57] where the two-dimensional electrodynamics under consideration was purely electric. A purely electric system can be obtained by reducing the third dimension ($\mu=2$) instead of the time coordinate. However, the resulting electrical model would still reside in Minkowskian spacetime. Such a model could not be mapped in a gravity model in an easy way as in the Euclidean case. This is the main reason why we perform dimensional reduction along the time direction rather than a spatial one. See [56, 57] for further details.
The static field equations can be obtained directly from the field equations (2.2) or from the minimization of the action (2.11). The equations for the scalar and vector fields are
@@
where $G$ and $C=C_{i}dx^{i}$ are integration constants that may, without loss of generality, be set to zero. These are massive Laplace equations whose solutions, in polar coordinates, are expressed in terms of the modified Bessel functions of the first and second kinds (depending on $mr$, with $r$ being the radial coordinate), and periodic functions in the angular coordinate.
It is also useful, for future purposes, to note that after the dimensional reduction, the gauge transformations (2.6) decouple to
@@
i.e., the scalar field becomes invariant under gauge transformations, while the magnetic vector potential remains as a gauge field.
As an extra and final step, we rewrite the action (2.11) with the aid of an auxiliary field (see [19, 56, 57]),
$S_{3qed}=\int\left\{-\frac{1}{\mathrm{e}^{2}}\left(d\Phi\ast d\Phi+\Theta\ast F-\frac{1}{2}\Theta\ast\Theta\right)+\kappa\left[2\Phi F-d\left(\Phi A\right)\right]\right\}\;.$ (2.16)
The field equation for the auxiliary field $\Theta$ trivially reads $\Theta=F$, which reproduces the action (2.11) again. Naturally, the introduction of auxiliary fields in this manner is also valid at the quantum level [68].
Finally, for completeness, the canonical dimensions and form ranks of the fields and parameters appearing in the action (2.16) are displayed in Table 1.
Fields AA $\Theta$ $\Phi$ $\mathrm{e}$ $\kappa$
Dimension 11 22 11 $1/2$ 0
Rank 11 22 0 0 0